OpenStudy (faiqraees):

Find the integral of \[\large\rm sin^{-1} x \ \] using the fact that \[\large\rm \frac{d}{dx}\sqrt{1-x^2} = \frac{-x}{\sqrt{1-x^2}} \]

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (faiqraees):

My approach was to find \[\large\rm ∫∫(-\frac{1}{x} \times \large\rm \frac{-x}{\sqrt{1-x^2}})\] Is this approach correct?

OpenStudy (phi):

Why a double integral?

OpenStudy (faiqraees):

Because we have to integrate arcsin x. So if we double integrate its derivative we can find its integral

OpenStudy (faiqraees):

But the problem is by using differentiation by parts I get a weird integral \[\large\rm ∫\frac{\sqrt{1-x^2}}{x^2}\]

OpenStudy (faiqraees):

Should I explain how I got to that double integral?

OpenStudy (phi):

when doing these problems, it's helpful to know the answer then you can work backwards.

OpenStudy (faiqraees):

Actually there are limits on the integral of 1 and 0. So the answer is in decimal notation

OpenStudy (phi):

ok. how do you get the double integral ?

OpenStudy (irishboy123):

they're setting you up for IBP \(\int \sin^{-1} x dx\) \(= \int (x)' \sin^{-1} x dx\) \(=x \sin^{-1} x - \int x (\sin^{-1} x)' dx\) \(=x \sin^{-1} x - \int x * \dfrac{1}{\sqrt{1-x^2}} dx\) and that second term is what you have been given

OpenStudy (faiqraees):

We are aware that d/dx arcsin x = 1/√1-x² Which means if I ∫∫1/√1-x² = ∫arcsin x And 1/√1-x² = -1/x * -x/√1-x²

OpenStudy (faiqraees):

@irishboy123 Oh got it thank you.

OpenStudy (phi):

I'm thinking you do \[ \int \sin^{-1} x \ dx = x \sin^{-1}- \int \frac{d}{dx} \sin^{-1 }x \ dx\]

OpenStudy (faiqraees):

shouldn't there be an 'x' in the last integral @phi

OpenStudy (phi):

yes. But Irish posted it correctly.